# Misspecification uncertainties in near-deterministic regression

### 02 Feb 2024 v1

@@ I Introduction
Surrogate models are widely used across science and engineering, notably to efficiently approximate the action of computationally intensive simulation engines  (Alizadeh et al., 2020; Deringer et al., 2019; Lapointe et al., 2020; Nyshadham et al., 2019; Montes de Oca Zapiain et al., 2022; Bonatti et al., 2022). Domain expertise is leveraged to design optimal features and model architecture before parameter inference is carried out based on training data generated by the engines. In this paper, we consider the common scenario where the learning problems share three key characteristics:
  • simulation engines are near deterministic: simulation outputs have vanishing aleatoric uncertainty. For example, the atomic energy $V({\bf X})$ is a deterministic function of atomic positions ${\bf X}$.
@@
[…]
Misspecification affects both finite capacity models and deep learning approaches with finite training resources  (Lahlou et al., 2021; Psaros et al., 2023). A misspecification-aware learning scheme should minimize the cross entropy between predicted and observed data distributions, known as the generalization error or predictive risk (see (Alquier, 2021; Morningstar et al., 2022)), but this loss is typically numerically intractable for high dimensional models.
Bayesian learning schemes instead minimize the expected loss (log likelihood), an upper bound to the generalization error, which is known to ignore misspecification. Epistemic uncertainties are therefore typically severely underestimated and can be shown to vanish in the underparametrized limit, with broad implications for surrogate model selection, uncertainty quantification and error propagation, as noted by multiple groups (Masegosa, 2020; Kato et al., 2022; Lotfi et al., 2022; Lahlou et al., 2021; Psaros et al., 2023; Imbalzano et al., 2021). We analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e. those typically used to approximate simulation engines in science and engineering. The deterministic limit is atypical in Bayesian analysis, as log likelihoods are typically required to be bounded and smooth, requiring a finite aleatoric uncertainty. Here, we take the deterministic limit to derive conditions that any parameter distribution must obey to avoid a divergent generalization error, then derive a simple ensemble ansatz which admits a robust learning scheme. Our main contributions are:
@@
  • For linear models, our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer. This is our main result, an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
  • We compare our ansatz to direct numerical minimization of the generalization error, showing how our scheme provides robust bounds on test errors. This is confirmed in numerical experiments on challenging datasets from atomic machine learning.
(Germain et al., 2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis, which bounds true expectations by empirical likelihood expectations through concentration inequalities. This provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization. In particular, minimizing the PAC-Bayes bound for the expected loss gives the familiar relation between posterior, prior and likelihood from Bayesian inference.
@@
(Masegosa, 2020) derived second order PAC-Bayes bounds for the generalization error, minimized through a variational or ensemble ansatz. (Morningstar et al., 2022) developed estimators to assess disagreement between the generalization error and expected loss, deriving specialized PAC bounds for theoretical guarantees. (Lahlou et al., 2021) considered misspecification of neural networks, jointly training a minimum loss surrogate and an independent predictor of misspecification error. (Lotfi et al., 2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error.
As misspecification has received only recent attention in the literature, various definitions for the distinction between epistemic and misspecification uncertainties exist (Kato et al., 2022). Here, misspecification is defined as non-aleatoric uncertainties which survive in the underparametrized limit $N/P\to\infty$, where PAC-Bayes bounds for the expected loss become tight. Our epistemic uncertainties thus coincide with standard estimates from loss minimization.
@@ II.1 Deterministic Simulation Engines
A simulation engine ${\mathcal{E}}:{\mathcal{X}}\to{\mathcal{Y}}$ takes input ${\mathbf{X}}\in{\mathcal{X}}$ and returns ${\mathbf{Y}}\in{\mathcal{Y}}$. We denote the dimension of ${\mathcal{Y}}$ and ${\mathcal{X}}$ by $Y$ and $X$ respectively. We consider the regime where the aleatoric uncertainty of ${\mathcal{E}}$ is weak and can be described as a Gaussian distribution with mean $\langle{\mathbf{Y}}|{\mathbf{X}}\rangle\equiv{\mathcal{E}}({\mathbf{X}})$ and covariance $\langle{\mathbf{Y}}\otimes{\mathbf{Y}}|{\mathbf{X}}\rangle-{\mathcal{E}}({\mathbf{X}})\otimes{\mathcal{E}}({\mathbf{X}})\equiv{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$, where $\|{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})\|=\mathcal{O}(\epsilon^{2Y})$, and $\epsilon$ is assumed small. The true conditional output distribution thus reads
@@
We assume homoskedatiscity ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})={\bm{\Sigma}}_{\mathcal{Y}}$ of this weak aleatoric error to investigate the deterministic, underparametrised limit $\epsilon\to 0$, $N/P\to\infty$ for $\epsilon=\mathcal{O}(\sqrt{P/N})$. With weights $\sum_{i=1}^{N}{\rm w}_{i}=1$, the training configurations ${\mathcal{D}}_{N}=\{{\mathbf{X}}_{i}\}_{i=1}^{N}$ form an input distribution
@@
The non-parametric estimate of $\rho_{N}$ will suffer from a curse of dimensionality (Tsybakov, 2008), i.e. the worst case error scales as $N^{1/(1+{\rm Dim}({\mathbf{X}}))}$$\simeq$1. However, learning only requires that we can bound expectations over independent identically distributed (i.i.d.) samples ${\mathbf{X}},{\mathbf{Y}}\in{\mathcal{D}}$
@@ II.2 Deterministic Surrogate Models and the Generalization Error
Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of the quasi-deterministic simulation engine ${\mathcal{E}}$. This has implications for model selection criteria, as we discuss in appendix A.
The generalization error ${\mathcal{G}}[\pi]$ is simply the cross entropy of predicted to observed distributions. The $\rho_{N}({\mathbf{X}})$ terms cancel inside the log and, omiting the constant ${\mathcal{G}}_{\mathcal{E}}\equiv\left\langle\ln\left|\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\right|\right\rangle_{\mathcal{D}}$, we have
${\mathcal{G}}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{D}},$ (6)
PAC-Bayes analysis (Alquier, 2021) provides concentration inequalities (Hoeffding, 1994) that bound expectations (3) over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$. These require $f({\mathbf{Y}},{\mathbf{X}})$ to be bounded (i.e. $\epsilon>0$) and expressed as an expectation over $\pi({\bm{\Theta}})$, i.e. $f({\mathbf{Y}},{\mathbf{X}})=\int_{\mathcal{P}}f({\mathbf{Y}},{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$. As ${\mathcal{G}}[\pi]$ is not in this form, PAC concentration inequalities can only be applied to some upper bound for ${\mathcal{G}}[\pi]$ which is an expectation over $\pi({\bm{\Theta}})$.
@@ II.3 The expected loss and PAC estimation
The bound used in Bayesian schemes is for the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood. The Jensen inequality for convex functions $f(\langle x\rangle)\leq\langle f(x)\rangle$ gives, with $f(x)=-\ln|x|$,
$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$ (7)
The empirical expected loss ^$\hat{{\mathcal{L}}}_{N}[\pi]$ over $\rho_{N}({\mathbf{X}})$ can estimate the true expected loss ${\mathcal{L}}[\pi]$ through the PAC-Bayes concentration inequality, as shown by (Germain et al., 2016)
${\mathcal{L}}[\pi]\leq\hat{{\mathcal{L}}}_{N}[\pi]$ $\equiv\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$
$+\frac{1}{N}\int_{\mathcal{P}}\ln\left|\frac{\pi({\bm{\Theta}})}{\pi_{0}({\bm{\Theta}})}\right|\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}+\frac{C_{0}}{N}$ (8)
where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})|\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, the distribution $\pi_{0}({\bm{\Theta}})$ is to some prior and $C_{0}$ is independent of $\pi$ but depends on $\epsilon$ and $\pi_{0}$. A well chosen prior thus tightens the PAC bound. Minimization of ^$\hat{{\mathcal{L}}}_{N}[\pi]$ yields the posterior familiar from Bayesian inference, as shown by  (Germain et al., 2016), but with an additional term $\lambda$ that ensures normalization-
@@
Appendix A discusses this bound in the context of Bayesian model selection for deterministic models. Using Laplace’s method as $N/P\to\infty$ gives $\hat{L}_{N}({\bm{\Theta}})\to L({\bm{\Theta}})$, with
@@
where $\langle\dots\rangle_{\mathcal{X}}\equiv\int\dots\rho_{\mathcal{D}}({\mathbf{X}}){\rm d}{\mathbf{X}}$. For linear problems, ${\bm{\Theta}}^{*}_{\mathcal{L}}$ is a generalized least squares solution, with the minimizing set a connected region in $\mathbb{R}^{P}$. In the underparametrized limit $N/P\to\infty$, expected loss is thus minimized by the sharp distribution $\pi_{\mathcal{L}}({\bm{\Theta}})\to\delta({\bm{\Theta}}-{\bm{\Theta}}^{*}_{\mathcal{L}})$, giving vanishing parameter uncertainty. The generalization error of the minimium loss solution clearly diverges as $1/\epsilon^{2}$:
$\lim_{N/P\to\infty}{\mathcal{G}}[\pi_{\mathcal{L}}]=\frac{1}{2}{\rm Tr}({\bm{\Sigma}^{*}_{\mathcal{L}}}{\bm{\Sigma}}_{\mathcal{Y}}^{-1})=\mathcal{O}(1/\epsilon^{2}),$ (12)
where ${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]\otimes[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$. As noted by (Masegosa, 2020), it is clear that the maximum likelihood parameter distribution $\pi_{\mathcal{L}}({\bm{\Theta}})$ is not a minimizer of the generalization error under misspecification.

### 09 Apr 2024 v2

@@ I Introduction
Surrogate models are widely used across science and engineering, notably to efficiently approximate the action of computationally intensive simulation engines  (Alizadeh et al., 2020; Deringer et al., 2019; Lapointe et al., 2020; Nyshadham et al., 2019; Montes de Oca Zapiain et al., 2022; Bonatti et al., 2022; Kudela and Matousek, 2022). Domain expertise is leveraged to design optimal features and model architecture before parameter inference is carried out based on training data generated by the engines. In this paper, we consider the common scenario where the learning problems share three key characteristics:
  • Simulation engines are near deterministic: simulation outputs have vanishing aleatoric uncertainty, e.g. the atomic energy $V({\bf X})$ is a deterministic function of atomic positions ${\bf X}$.
@@
  • Surrogate models are underparametrized: large quantities of training data are available, such that the standard posterior from Bayesian inference gives vanishing parameter uncertainties. This typically occurs in the limit $P/N\ll 1$ for finite capacity models.
Misspecification affects both finite capacity models and deep learning approaches with finite training resources  (Lahlou et al., 2021; Psaros et al., 2023). A misspecification-aware learning scheme should minimize the cross entropy between predicted and observed data distributions, known as the generalization error, or population risk (see (Alquier, 2021; Morningstar et al., 2022)). However, this is typically not tractable for learning, both due to numerical instabilities and the lack of theoretical bounds for convergence.
Bayesian learning schemes instead minimize the expected loss (log likelihood), an upper bound to the generalization error, which is known to ignore misspecification but admits a robust learning scheme. Epistemic uncertainties are thus severely underestimated, vanishing in the underparametrized limit, with broad implications for surrogate model selection, uncertainty quantification and error propagation, as noted by multiple groups(Masegosa, 2020; Kato et al., 2022; Lotfi et al., 2022; Lahlou et al., 2021; Psaros et al., 2023; Imbalzano et al., 2021).
We analyze the generalization error of deterministic, misspecified and under-parametrized surrogate models, i.e. those typically used to approximate simulation engines in science and engineering. Our focus is the underparametrized regime, meaning we do not focus on epistemic errors important for finite $N$. As posterior distributions become unbounded in the deterministic limit, this is only taken as a theoretical device to derive conditions that any posterior must obey to avoid a divergent generalization error. We then design an ensemble ansatz that respects this condition and admits a robust learning scheme at finite $N$, which reduces to a constrained loss minimization problem for large $N/P$. Our main contributions are:
@@
  • For linear models, the variational minimum of our ansatz can be efficiently evaluated via rank-one updates to a leverage-weighted loss minimizer. This is our main result, an efficient scheme to incorporate misspecification into linear regression with minimal computational overhead.
  • We compare our variational minimum to numerial optimization of our ensemble ansatz and show how our scheme provides robust bounds on test errors. This is confirmed in numerical experiments on challenging datasets from atomic machine learning.
(Germain et al., 2016) established an important connection between Bayesian inference and probability approximately correct (PAC-Bayes) analysis for regression problems with unbounded losses, which bounds true expectations by empirical likelihood expectations through concentration inequalities. This provides a robust learning scheme without test-train splitting, and gives a rationalization for Bayesian inference as a form of regularization. In particular, minimizing the PAC-Bayes bound for the expected loss gives the familiar relation between posterior, prior and likelihood from Bayesian inference.
@@
(Masegosa, 2020) derived second order PAC-Bayes bounds for the generalization error, minimized through a variational or ensemble ansatz. (Morningstar et al., 2022) developed estimators to assess disagreement between the generalization error and expected loss, deriving specialized PAC bounds for theoretical guarantees. (Lahlou et al., 2021) considered misspecification of neural networks, jointly training a minimum loss surrogate and an independent predictor of misspecification error. (Lotfi et al., 2022) considered misspecification in the context of deep model selection, deriving conditional likelihoods better aligned with the generalization error. As various definitions for the distinction between epistemic and misspecification uncertainties exist (Kato et al., 2022), we define misspecification as non-aleatoric uncertainties which survive in the underparametrized limit, where the minimizer of PAC-Bayes bounds for the expected loss, which coincides with the posterior from Bayesian inference, predict vanishing parameter uncertainties. This definition is made more precise in section II.3. Our epistemic uncertainties thus coincide with standard estimates from loss minimization.
We summarize known results for the generalization error and expected loss, PAC-Bayes bounds for the regression loss, and the sub-optimality of loss minimization for misspecified problems in section II. Section III contains our central theoretical contribution, an analysis of the generalization error for near-deterministic surrogate models and a simple ensemble ansatz which can be efficiently evaluated for linear models. Section IV applies our ansatz to high-dimensional linear regression problems, using synthetic data and challenging applications in atomic machine learning.
@@ II.1 Deterministic Simulation Engines
A simulation engine ${\mathcal{E}}:{\mathcal{X}}\to{\mathcal{Y}}$ takes input ${\mathbf{X}}\in{\mathcal{X}}$ and returns ${\mathbf{Y}}\in{\mathcal{Y}}$. We consider the regime where the aleatoric uncertainty of ${\mathcal{E}}$ is weak and can be described as a Gaussian distribution with mean $\langle{\mathbf{Y}}|{\mathbf{X}}\rangle\equiv{\mathcal{E}}({\mathbf{X}})$ and covariance $\langle{\mathbf{Y}}{\mathbf{Y}}^{\top}|{\mathbf{X}}\rangle-{\mathcal{E}}({\mathbf{X}}){\mathcal{E}}^{\top}({\mathbf{X}})\equiv{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})$, where $\|{\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})\|=\mathcal{O}(\epsilon^{2Y})$, and $\epsilon$ is assumed small. The true conditional output distribution thus reads
@@
We assume homoskedatiscity ${\bm{\Sigma}}_{\mathcal{Y}}({\mathbf{X}})={\bm{\Sigma}}_{\mathcal{Y}}$ of this weak aleatoric error to investigate the deterministic, underparametrised limit $\epsilon\to 0$, $N/P\to\infty$ for $\epsilon=\mathcal{O}(\sqrt{P/N})$. With weights $\sum_{i=1}^{N}{\rm w}_{i}=1$, the training configurations ${\mathcal{D}}_{N}=\{{\mathbf{Y}}_{i},{\mathbf{X}}_{i}\}_{i=1}^{N}$ form an input distribution in ${\mathbf{X}}$ of
@@
Learning requires that we can bound expectations over independent identically distributed (i.i.d.) samples ${\mathbf{X}},{\mathbf{Y}}\in{\mathcal{D}}$
@@ II.2 Deterministic Surrogate Models and the Generalization Error
Unlike typical Bayesian schemes, we do not treat the covariance term ${\bm{\Sigma}}_{\mathcal{Y}}$ as fitting parameter but as an intrinsic property of the quasi-deterministic simulation engine ${\mathcal{E}}$. This has implications for model selection criteria, as we discuss in appendix B.
The generalization error ${\mathcal{G}}[\pi]$ is simply the cross entropy of the predicted distribution $\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$ to the observed distribution $\rho({\mathbf{Y}}|{\mathbf{X}})\rho_{N}({\mathbf{X}})$; omiting the constant term ${\mathcal{G}}_{\mathcal{E}}\equiv\left\langle\ln\left|\rho_{\mathcal{E}}({\mathbf{Y}}|{\mathbf{X}})\right|\right\rangle_{\mathcal{D}}$ we have
${\mathcal{G}}[\pi]=-\left\langle\ln\left|\int_{\mathcal{P}}\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}\right|\right\rangle_{\mathcal{D}}.$ (6)
@@ II.3 The expected loss and PAC estimation
PAC-Bayes analysis (Alquier, 2021) provides concentration inequalities (Hoeffding, 1994) that bound expectations (3) over ${\mathcal{D}}$ by empirical expectations over ${\mathcal{D}}_{N}$ with some probability $1-\xi$. Whilst originally designed for classification problems where the loss is bounded to $[0,1]$, the treatment can be extended to regression settings with unbounded losses, as discussed in (Germain et al., 2016) and appendix A. In this setting, available PAC-Bayes bounds are for $\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{\mathcal{D}}$, where $f({\mathbf{Y}},{\mathbf{X}})$ to be bounded (i.e. $\epsilon>0$) and can expressed as an expectation over $\pi({\bm{\Theta}})$, i.e. $\langle f({\mathbf{Y}},{\mathbf{X}})\rangle_{\mathcal{D}}=\int_{\mathcal{P}}\langle f({\mathbf{Y}},{\mathbf{X}},{\bm{\Theta}})\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}$. As ${\mathcal{G}}[\pi]$ is not in this form, the bound used in PAC-Bayesian schemes is for the expected loss ${\mathcal{L}}[\pi]\geq{\mathcal{G}}[\pi]$, often known as the negative log likelihood. The Jensen inequality for convex functions $f(\langle x\rangle)\leq\langle f(x)\rangle$ gives, with $f(x)=-\ln|x|$,
$\mathcal{G}[\pi]\leq{\mathcal{L}}[\pi]=-\int_{\mathcal{P}}\left\langle\ln\left|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\right|\right\rangle_{\mathcal{D}}\pi({\bm{\Theta}}){\rm d}{\bm{\Theta}}.$
As discussed by (Germain et al., 2016) the expected regression loss ${\mathcal{L}}[\pi]$ admits the PAC-Bayes upper bound, holding with probability $1-\xi$, of
${\mathcal{L}}[\pi]$$\leq\hat{{\mathcal{L}}}_{N}[\pi]+\frac{1}{N}\left(KL(\pi|\pi_{0})-\ln|\xi|\right)+C_{0}[\pi_{0}],$
^$\hat{{\mathcal{L}}}_{N}[\pi]$ $\equiv\int_{\mathcal{P}}\hat{L}_{N}({\bm{\Theta}}){\rm d}\pi({\bm{\Theta}}),$ (7)
where $\hat{L}_{N}({\bm{\Theta}})=-\langle\ln|\rho_{\mathcal{M}}({\mathbf{Y}}|{\mathbf{X}},{\bm{\Theta}})\rangle_{N}$ is the empirical loss over $\rho_{N}({\mathbf{X}})$, $KL(\pi|\pi_{0})=\int_{\mathcal{P}}\ln|\pi/\pi_{0}|{\rm d}\pi({\bm{\Theta}})$ is the cross entropy between $\pi$ and some prior $\pi_{0}({\bm{\Theta}})$ and $C_{0}[\pi_{0}]$ is constant with respect to $\pi$ and has a finite upper bound discussed in appendix A. The minimizer of the PAC-Bayes upper bound (7) is that of ^$\hat{{\mathcal{L}}}_{N}[\pi]+KL(\pi|\pi_{0})/N$, giving the posterior from Bayesian inference (Germain et al., 2016)-
@@
Appendix B discusses this bound in the context of Bayesian model selection. Using Laplace’s method (appendix C) as $N/P\to\infty$ gives $\hat{L}_{N}({\bm{\Theta}})\to L({\bm{\Theta}})$, with
@@
[…]
$\lim_{N/P\to\infty}{\mathcal{G}}[\pi_{\mathcal{L}}]=\frac{1}{2}{\rm Tr}({\bm{\Sigma}^{*}_{\mathcal{L}}}{\bm{\Sigma}}_{\mathcal{Y}}^{-1})=\mathcal{O}(1/\epsilon^{2}).$ (11)
${\bm{\Sigma}}^{*}_{\mathcal{L}}=\langle[{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})][{\mathcal{E}}({\mathbf{X}})-{\mathcal{M}}({\mathbf{X}},{\bm{\Theta}}^{*}_{\mathcal{L}})]^{\top}\rangle$ is the error covariance around ${\bm{\Theta}}^{*}_{\mathcal{L}}$ and we use ${\rm Tr}({\bm{\Sigma}}_{\mathcal{Y}})=\mathcal{O}(\epsilon^{2})$. As noted by (Masegosa, 2020), it is clear that the maximum likelihood parameter distribution $\pi_{\mathcal{L}}({\bm{\Theta}})$ is not a minimizer of the generalization error under misspecification.
We emphasize that the primary motivation of this section is to show that whilst the generalization error is the central quantity of interest for learning under misspecification, we typically only have access to an upper bound in the form of a PAC-Bayes concentration inequality (7) for the expected loss. Under model misspecification it is clear that minimization of this upper bound in sub-optimal, in that the generalization error diverges as $1/\epsilon^{2}$. In the next section we analyze the generalization error as $\epsilon\to 0$ to derive a condition for $\pi({\bm{\Theta}})$ that ensures a finite generalization error. We then design an ensemble ansatz which respects this condition and can be parametrized through a variational minimization of the PAC-Bayes upper bound (7).